# Detection of very high-energy gamma-ray emission from the radio galaxy M87 with LHAASO

### v1

@@ 4.2 Long-term low-state VHE emission
The LHAASO observations revealed a long-term VHE emission with a flux level of $\sim 10^{-12}~{}\rm erg~{}cm^{-2}~{}s^{-1}$. The intrinsic spectrum can be described by a power law, with a possible hardening at 20 TeV. Although for the episodes of rapid TeV flaring activity an origin in the sub-parsec scale jet and below is preferred, it could well be that larger scale structures (such as HST-1, the kpc-scale jet and the knot A) contribute to the overall, low-state VHE emission. The nuclear environment of M87 does not show evidence of significant sources of soft photons (e.g., from the torus or star formation sites) and therefore the inverse Compton emission is likely dominated by the synchrotron self-Compton (SSC) component from the same population of relativistic electrons that produce the synchrotron photons. However, a simple one-zone synchrotron and SSC model cannot describe the observed SED easily (Tavecchio & Ghisellini, 2008). The reason for that is the following, in order to have the synchrotron component peaking in the infrared (IR) region and the SSC component close to the TeV band (as required by the relatively flat TeV spectra), one has to assume an unreasonably large Doppler factor, $\delta\gg 10$. The hard TeV spectrum also challenges the one-zone SSC model, since the Klein-Nishina effect leads to a spectral steepening towards the highest energy end. Indeed, previous detailed modelings of the low-state emission have found that the SSC model provides a good fit to the SED up to GeV gamma-rays, but it underestimates the VHE emission (e.g. Ait Benkhali et al., 2019; Alfaro et al., 2022).
@@
To overcome this problem, one should “decouple” the synchrotron and IC components, assuming that the two components are produced in two different regions, as in the “decelerating jet” model (Georganopoulos et al., 2005) or in the “spine-layer” model (Tavecchio & Ghisellini, 2008). Other possibility includes several randomly oriented active regions resulting from reconnection events in the jet (the so-called “jets-in-the-jet” scenario Giannios et al., 2009, 2010). An alternative model is a lepto-hadronic emission model (Reimer et al., 2004; Fraija & Marinelli, 2016; Boughelilba et al., 2022), in which three spectral components are involved: a synchrotron-dominated component from radio to X-rays, an inverse Compton-dominated component from X-rays to GeV gamma-rays, and a photohadronic-dominated component in the VHE band.
@@
4.3 Constraining the upper bound of the EBL intensity with the VHE data
@@
The EBL is not well established with large uncertainties in the flux. Our results allow us to put a constraint on the upper bound of the EBL intensity, assuming that the intrinsic spectrum is described by a single power law extending to the maximum energy at 20 TeV. We adopt a model for the observed spectrum that incorporates a power-law multiplied by the scaled EBL absorption: $dN/dE=N_{0}(E/{\rm 3TeV})^{-\gamma}\times\textrm{exp}(-f_{\rm EBL}\tau(E))$, where $\tau(E)$ is the absorption depth due to EBL for a gamma-ray photon with energy $E$ and $f_{\rm EBL}$ is the scaling factor. We use $\tau(E)$ from the Franceschini et al. (2008) as reference values. Since the absorption for low energy photons below 10 TeV is negligible ($\tau\leq 0.1$), the relevant wavelength range [^2] This wavelength range is obtained from the peak of the pair production cross section at the wavelength for a specific gamma-ray photon of energy E: $\lambda_{\rm max}\simeq 1.24(E/{\rm TeV})~{}\mu m$. for EBL constraint is $\lambda\gtrsim 10~{}\mu m$. To obtain the upper limit of $f_{\rm EBL}$, we estimate the likelihood values by varying the parameter $f_{\rm EBL}$ in fitting the spectral data observed by LHAASO. The result is shown in Figure 7. We find that the upper limit of $f_{\rm EBL}$ at a 95% confidence level is about 1.6. This means that the upper bound of the EBL intensity is a factor of 1.6 larger than the intensity given by Franceschini et al. (2008) for the wavelength range $\lambda\gtrsim 10~{}\mu m$. We also study the constraint on the dominguez-upper model that reflects maximum EBL flux (Domínguez et al., 2011). We find that the upper limit of $f_{\rm EBL}$ at a 95% confidence level is only 1.09. This means that the upper limit allowed by our TeV data is very close to the maximum EBL flux model.
4.4 Influence of the possible internal absorption on the spectrum
@@
TeV gamma-rays could be absorbed by the radiation fields in the central region of M87. The observed infrared luminosity of the nucleus of M87 is about $10^{40}-10^{41}{\rm erg~{}s^{-1}}$ (Perlman et al., 2001; Whysong & Antonucci, 2004). Although there are no direct measurements of the size of the infrared source, observations in the microwave band at 43 GHz suggest that its size at millimeter wavelengths is limited by $5\times 10^{16}{\rm cm}$, corresponding to 25 $R_{s}$ for a BH mass of $\sim 6\times 10^{9}M_{\odot}$. Neronov & Aharonian (2007) calculated the optical depth for gamma-rays produced by the infrared source assuming that the size of the infrared source is comparable to that of the microwave source, finding that the nucleus of M87 is effectively transparent for gamma-rays up to an energy of 10 TeV. Brodatzki et al. (2011) consider the temperature distribution of the disk photons of M87 and calculate the optical depth of gamma-rays for different inclination angle $\theta$ between the jet and our line of sight. They find that, for $\theta\sim 20^{\circ}$, the optical depth exceeds unity only for photons with energy above 20 TeV.
@@
We perform a re-analysis of the VHE spectra of M87 assuming internal absorption based on the above two models, i.e., model 1 for Neronov & Aharonian (2007) and model 2 for Brodatzki et al. (2011). The intrinsic VHE spectra of M87 during the full-time period and the flare period after considering the internal absorption are shown in Figure 8. It is found that the spectra become harder with a spectral index being $1.86\pm 0.21$ for model 1 and $2.06\pm 0.20$ for model 2 during the full-time period. During the flare period, the spectral index is $2.26\pm 0.31$ for model 1 and $2.41\pm 0.30$ for model 2. The excess at 20 TeV becomes more pronounced in the VHE spectra during the full-time period. Note that, in the calculation, we have assumed that the emitting region of the VHE emission during the low state is also close the AGN core, similar to the case during the flare state.
5 Summary
@@ Appendix A LHAASO Data Reduction
The dataset utilized in this analysis was mainly sourced from the WCDA at LHAASO. For this study, we restricted our analysis to the events with a zenith angle of less than 50 degrees to ensure high-quality reconstruction data. The selected events are subsequently categorized into seven groups based on the effective number of triggered PMT units, denoted as $N_{\rm hit}$, with ranges [60-100), [100-200), [200-300), [300-500), [500-700), [700-1000) and [1000, 2000]. Additionally, a gamma/proton separation parameter $\mathcal{P}_{\rm{ciness}}$ is employed to effectively identify gamma-like events. $\mathcal{P}_{\rm{ciness}}$ cut of seven $N_{\textrm{hit}}$ segments are 1.02, 0.90, 0.88, 0.88, 0.84, 0.84 and 0.84. Ultimately, the effective live-time for observations of M87 was recorded as 1026 days, with an approximate count of $9.61\times 10^{9}$ gamma-like events. The event maps are created as histograms of the arrival direction of the reconstructed events, after converting the arrival direction of the shower from local coordinates to equatorial coordinates. The background maps are computed using a direct integration method (Fleysher et al., 2004). For this analysis, a sliding time window of 10 hours was used to estimate the detector acceptance, with an integration time set to 4 hour to estimate the relative back-ground. Events within the region of the Galactic plane and LHAASO 1st catalog gamma-ray sources (with a spatial angle less than 5 degrees) were excluded from the background estimation. The excess map is then obtained by subtracting the background map from the event data map.
For the spectral fitting, we assume a simple power-law function for the intrinsic spectrum. The photon attenuation derived from the three EBL models (e.g., the kneiske model (Kneiske & Dole, 2010), franceschini model (Franceschini et al., 2008), and dominguez-upper model (Domínguez et al., 2011)) as a function of energy at the distance z = 0.0042 is applied to correct the detector effective area, via re-weighting every simulated event with the survival probability $\mathcal{P}(E)=e^{-\tau(E)}$ from the EBL attenuation, where $\tau(E)$ is the optical depth due to the EBL absorption at the photon energy $E$. The results are summarized in Table 1, indicating that the impact of different EBL models on the maximum energy is not significant.

### v3

@@ IV.2 Long-term low-state VHE emission
The LHAASO observations revealed a long-term VHE emission with a flux level of $\sim 10^{-12}~\rm erg~cm^{-2}~s^{-1}$. The intrinsic spectrum can be described by a power law, with a possible hardening at 20 TeV. Although for the episodes of rapid TeV flaring activity an origin in the sub-parsec scale jet and below is preferred, it could well be that larger scale structures (such as HST-1, the kpc-scale jet and the knot A) contribute to the overall, low-state VHE emission. The nuclear environment of M87 does not show evidence of significant sources of soft photons (e.g., from the torus or star formation sites) and therefore the inverse Compton emission is likely dominated by the synchrotron self-Compton (SSC) component from the same population of relativistic electrons that produce the synchrotron photons. However, a simple one-zone synchrotron and SSC model cannot describe the observed SED easily (2008MNRAS.385L..98T). The reason for that is the following, in order to have the synchrotron component peaking in the infrared (IR) region and the SSC component close to the TeV band (as required by the relatively flat TeV spectra), one has to assume an unreasonably large Doppler factor, $\delta\gg 10$. The hard TeV spectrum also challenges the one-zone SSC model, since the Klein-Nishina effect leads to a spectral steepening towards the highest energy end. Indeed, previous detailed modelings of the low-state emission have found that the SSC model provides a good fit to the SED up to GeV gamma-rays, but it underestimates the VHE emission (e.g. 2019A&A...623A...2A; 2022ApJ...934..158A).
@@
To overcome this problem, one should “decouple” the synchrotron and IC components, assuming that the two components are produced in two different regions, as in the “decelerating jet” model (2005ApJ...634L..33G) or in the “spine-layer” model (2008MNRAS.385L..98T). Other possibility includes several randomly oriented active regions resulting from reconnection events in the jet (the so-called “jets-in-the-jet” scenario 2009MNRAS.395L..29G; 2010MNRAS.402.1649G). An alternative model is a lepto-hadronic emission model (2004A&A...419...89R; 2016ApJ...830...81F; 2022ApJ...938...79B), in which three spectral components are involved: a synchrotron-dominated component from radio to X-rays, an inverse Compton-dominated component from X-rays to GeV gamma-rays, and a photohadronic-dominated component in the VHE band.
@@
IV.3 Constraining the upper bound of the EBL intensity with the VHE data
@@
The EBL is not well established with large uncertainties in the flux. Our results allow us to put a constraint on the upper bound of the EBL intensity, assuming that the intrinsic spectrum is described by a single power law extending to the maximum energy at 20 TeV. We adopt a model for the observed spectrum that incorporates a power-law multiplied by the scaled EBL absorption: $dN/dE=N_{0}(E/{\rm 3TeV})^{-\gamma}\times\textrm{exp}(-f_{\rm EBL}\tau(E))$, where $\tau(E)$ is the absorption depth due to EBL for a gamma-ray photon with energy $E$ and $f_{\rm EBL}$ is the scaling factor. We use $\tau(E)$ from the 2008A&A...487..837F as reference values. Since the absorption for low energy photons below 10 TeV is negligible ($\tau\leq 0.1$), the relevant wavelength range [^2] This wavelength range is obtained from the peak of the pair production cross section at the wavelength for a specific gamma-ray photon of energy E: $\lambda_{\rm max}\simeq 1.24(E/{\rm TeV})~\mu m$. for EBL constraint is $\lambda\gtrsim 10~\mu m$. To obtain the upper limit of $f_{\rm EBL}$, we estimate the likelihood values by varying the parameter $f_{\rm EBL}$ in fitting the spectral data observed by LHAASO. The result is shown in Figure 7. We find that the upper limit of $f_{\rm EBL}$ at a 95% confidence level is about 1.6. This means that the upper bound of the EBL intensity is a factor of 1.6 larger than the intensity given by 2008A&A...487..837F for the wavelength range $\lambda\gtrsim 10~\mu m$. We also study the constraint on the dominguez-upper model that reflects maximum EBL flux (2011MNRAS.410.2556D). We find that the upper limit of $f_{\rm EBL}$ at a 95% confidence level is only 1.09. This means that the upper limit allowed by our TeV data is very close to the maximum EBL flux model.
IV.4 Influence of the possible internal absorption on the spectrum
@@
TeV gamma-rays could be absorbed by the radiation fields in the central region of M87. The observed infrared luminosity of the nucleus of M87 is about $10^{40}-10^{41}{\rm erg~s^{-1}}$ (2001ApJ...551..206P; 2004ApJ...602..116W). Although there are no direct measurements of the size of the infrared source, observations in the microwave band at 43 GHz suggest that its size at millimeter wavelengths is limited by $5\times 10^{16}{\rm cm}$, corresponding to 25 $R_{s}$ for a BH mass of $\sim 6\times 10^{9}M_{\odot}$. 2007ApJ...671...85N calculated the optical depth for gamma-rays produced by the infrared source assuming that the size of the infrared source is comparable to that of the microwave source, finding that the nucleus of M87 is effectively transparent for gamma-rays up to an energy of 10 TeV. 2011ApJ...736...98B consider the temperature distribution of the disk photons of M87 and calculate the optical depth of gamma-rays for different inclination angle $\theta$ between the jet and our line of sight. They find that, for $\theta\sim 20^{\circ}$, the optical depth exceeds unity only for photons with energy above 20 TeV.
@@
We perform a re-analysis of the VHE spectra of M87 assuming internal absorption based on the above two models, i.e., model 1 for 2007ApJ...671...85N and model 2 for 2011ApJ...736...98B. The intrinsic VHE spectra of M87 during the full-time period and the flare period after considering the internal absorption are shown in Figure 8. It is found that the spectra become harder with a spectral index being $1.86\pm 0.21$ for model 1 and $2.06\pm 0.20$ for model 2 during the full-time period. During the flare period, the spectral index is $2.26\pm 0.31$ for model 1 and $2.41\pm 0.30$ for model 2. The excess at 20 TeV becomes more pronounced in the VHE spectra during the full-time period. Note that, in the calculation, we have assumed that the emitting region of the VHE emission during the low state is also close the AGN core, similar to the case during the flare state.
V Summary
@@
Figure 1: Significance maps of $2^{\circ}\times 2^{\circ}$ region around M87 by LHAASO-WCDA for full-time period (left panel), flare state (middle panel) and low state (right panel), respectively. M87 is marked by cyan open cross. The gray circle indicates the 95% position error of the source. The white circle at the bottom-left corner shows the size of the LHAASO PSF (68% containment).
Figure 2: Observed VHE spectra of M87 by LHAASO during the full-time period assuming the PLEC function and LP function, respectively. The solid lines indicate the best-fitting results and the blue shaded regions indicate the $1\sigma$ statistical error. Left: The PLEC function is adopted to fit the observational data. Right: The LP function is adopted to fit the observational data.
Figure 3: VHE spectra of M87 during the full-time period (left panel), the flare state (middle panel) and the low state (right panel), assuming a single power-law model for the intrinsic spectra and a EBL absorption model of 2008A&A...487..837F. The gray and blue shaded regions indicate $1\sigma$ statistical error for the intrinsic spectra and PLxEBL spectra, respectively.
Figure 4: Light curves of LHAASO-WCDA (top panels) and Fermi-LAT (bottom panels). Left panel: long-term light curves (three-month bin) of LHAASO-WCDA and Fermi-LAT for M87. Right panel: zoom-in light curves around the flare period. The light curve of LHAASO-WCDA is obtained with two-day bin and that of Fermi-LAT is obtained with 8-day bin. $1\sigma$ upper limit is reported when $\rm TS<9$ for the LAT light curve data. The LHAASO-WCDA flare period is shown as the gray vertical shaded region. The blue line in the top panel indicates the bayesian block and the gray lines show the long-term average flux.
Figure 5: Fitting of the one-day bin light curve (range from MJD 59593 to MJD 59629) of the VHE flare measured by LHAASO-WCDA with a two-sided exponential function (the red line). Fitting parameters are $F_{0}=(5.62\pm 1.26)\times 10^{-12}~\rm TeV^{-1}~cm^{-2}~s^{-1}$, $t_{0}=\rm MJD~59610\pm 0.58$, $\tau_{d}^{\rm rise}=1.05\pm 0.49$ days and $\tau_{d}^{\rm decay}=2.17\pm 0.58$ days.
Figure 6: Gamma-ray SED of M87 for the low state (magenta data points) and flare state (blue data points), respectively. The flare state encompasses data from MJD 59607 to MJD 59615. The VHE data of LHAASO show the intrinsic spectra, and the shaded regions indicate $1\sigma$ statistical error for the fitting of the intrinsic spectra. The gray data points represent the observation of H.E.S.S. in the low state (2023A&A...675A.138H).
Table 1: Best-fit parameters of the intrinsic spectrum for the full-time LHAASO data based on three EBL models.
EBL model $N_{0}@{\rm 3TeV}$ $\gamma$ $E_{\gamma,\rm{max}}$ TS $\chi^{2}/{\rm ndf}$
($\rm{10^{-13}~TeV^{-1}~cm^{-2}~s^{-1}}$) (TeV)
dominguez-upper $0.323\pm 0.055$ $2.25\pm 0.18$ 18.4 87.5 3.47/4
franceschini $0.327\pm 0.050$ $2.37\pm 0.14$ 19.4 87.4 3.00/4
kneiske $0.330\pm 0.044$ $2.56\pm 0.12$ 20.6 84.4 3.97/4
Table 2: Best-fit spectral parameters of the LHAASO data for different time intervals.
Time interval State TS $N_{0}@{\rm 3TeV}$ $\gamma$ Flux
(MJD) ($\rm{TeV^{-1}~cm^{-2}~s^{-1}}$) ($\rm{pho~cm^{-2}~s^{-1}}$)
[…]
59616–60385
Table 3: Best-fit spectral parameters of the Fermi-LAT data.
Time interval TS $N_{0}$ $\Gamma_{LAT}$ Flux
(MJD) ($\rm{ph~cm^{-2}~s^{-1}}$) ($\rm{pho~cm^{-2}~s^{-1}}$)
[…]
Figure 7: The likelihood profile on fitting the spectral data observed by LHAASO for a varying scale factor $f_{\rm EBL}$ based on the EBL model of 2008A&A...487..837F (left panel) and the EBL model of 2011MNRAS.410.2556D (right panel). The grey dashed line represents $2({\rm log}L-{\rm log}L_{0})=2.71$, which corresponds to a 95% confidence level (where $L$ is the likelihood for different scale factors $f_{\rm EBL}$ and $L_{0}$ is the maximum likelihood ).
Figure 8: Intrinsic VHE spectra of M87 during the full-time period (left panel) and the flare period (right panel) after considering the internal absorption modeled by 2007ApJ...671...85N (model 1) and Brodatzki_2011 (model 2).
@@ Appendix A LHAASO Data Reduction
The dataset utilized in this analysis was mainly sourced from the WCDA at LHAASO. For this study, we restricted our analysis to the events with a zenith angle of less than 50 degrees to ensure high-quality reconstruction data. The selected events are subsequently categorized into seven groups based on the effective number of triggered PMT units, denoted as $N_{\rm hit}$, with ranges [60-100), [100-200), [200-300), [300-500), [500-700), [700-1000) and [1000, 2000]. Additionally, a gamma/proton separation parameter $\mathcal{P}_{\rm{ciness}}$ is employed to effectively identify gamma-like events. $\mathcal{P}_{\rm{ciness}}$ cut of seven $N_{\textrm{hit}}$ segments are 1.02, 0.90, 0.88, 0.88, 0.84, 0.84 and 0.84. Ultimately, the effective live-time for observations of M87 was recorded as 1026 days, with an approximate count of $9.61\times 10^{9}$ gamma-like events. The event maps are created as histograms of the arrival direction of the reconstructed events, after converting the arrival direction of the shower from local coordinates to equatorial coordinates. The background maps are computed using a direct integration method (Fleysher_2004). For this analysis, a sliding time window of 10 hours was used to estimate the detector acceptance, with an integration time set to 4 hour to estimate the relative back-ground. Events within the region of the Galactic plane and LHAASO 1st catalog gamma-ray sources (with a spatial angle less than 5 degrees) were excluded from the background estimation. The excess map is then obtained by subtracting the background map from the event data map.
For the spectral fitting, we assume a simple power-law function for the intrinsic spectrum. The photon attenuation derived from the three EBL models (e.g., the kneiske model (2010A&A...515A..19K), franceschini model (2008A&A...487..837F), and dominguez-upper model (2011MNRAS.410.2556D)) as a function of energy at the distance z = 0.0042 is applied to correct the detector effective area, via re-weighting every simulated event with the survival probability $\mathcal{P}(E)=e^{-\tau(E)}$ from the EBL attenuation, where $\tau(E)$ is the optical depth due to the EBL absorption at the photon energy $E$. The results are summarized in Table 1, indicating that the impact of different EBL models on the maximum energy is not significant.